A coin bank contains 25 coins in nickels, dimes, and quarters. There are four times as many dimes as quarters. The value of the coins is $2.05. How many dimes are in the bank?
Answer by ewatrrr(24785)


Hi
bank contains 25 coins with four times as many dimes as quarters
Let x, 4x and (25-5x) represent quarters, dimes, nickels respectively
Questions states***Cents makes sense!
25x + 10*4x + 5(25-5x) = 205 CENTS
solving for x
65x + 125 - 25x = 205
40x = 80
x = 2, the number of quarters, dimes number 8 (15 nickels = 25-10)
CHECKING our Answer
25*2 + 10*8 + 5*15 = 205 or $2.05